OpenStudy (anonymous):

Determine if the given sequence is an arithmetic sequence. If it is, find the common difference, D. N: 1 2 3 4 f(n): 10 12 14 16 Arithmetic sequence: Common Difference:

OpenStudy (tkhunny):

Well, is it?

Directrix (directrix):

Arithmetic Sequence A sequence such as 1, 5, 9, 13, 17, ... which has a constant difference between terms. The finite sequence 10 12 14 16 has a constant difference of 2 between adjacent terms so it is an Arithmetic sequence. @PECKCE

OpenStudy (tkhunny):

N: 1 2 3 4 f(n): 10 12 14 16 Subtract successive values: 12 - 10 = 2 14 - 12 = 2 16 - 14 = 2 It's not overwhelming evidence, but it's a pretty good hint that it is an arithmetic sequence and the common difference is 2.

OpenStudy (anonymous):

That what I though just was unsure. thank you guys.